# Nash Equilibria in Games with Playerwise Concave Coupling Constraints: Existence and Computation

### v1

@@ 1 INTRODUCTION
Optimization problems involving multiple self-interested agents with coupled objectives frequently arise in applications across robotics, computer science, and economics. Game theory provides a natural framework for modelling such interactions. In many interactive tasks, players are restricted to optimizing over a set of admissible strategies that are influenced by the choices of others. For instance, players may be subject to satisfying safety requirements such as collision avoidance in multi-robot control (Gu et al.,, 2023), or operating within a shared budget, e.g. limited bandwidth in communication networks (Lasaulce and Tembine,, 2011). This more general game setting is captured by introducing coupling constraints, i.e., constraints that depend on the players’ joint strategy profile. In economic modeling, such constraint games have been applied to analyze equilibria under resource limitations and market constraints (Arrow and Debreu,, 1954).
@@
In the following, we distinguish between games with non-shared and shared coupling constraints. In the former, more general setting, each player may face distinct coupling constraints—a framework introduced by Debreu, (1952). Therein, constrained Nash equilibria, referred to as social equilibria, are shown to exist in games with compact, convex strategy spaces and playerwise quasi-concave utilities, provided that each player has a nonempty set of feasible response strategies, irrespective of the strategies chosen by others. This assumption is strong: in many practical scenarios, such as under shared safety constraints, violations caused by a subset of players may leave others unable to restore feasibility. However, without any additional assumptions, we give a simple example of a game with linear utilities and non-shared linear coupling constraints for which no constrained Nash equilibrium exists (see Appendix B).
This motivates studying the latter case: shared coupling constraints, where the set of admissible joint strategies is identical across players. In this setting, Rosen, (1965) showed that a constrained Nash equilibrium exists under the assumption of joint convexity of the feasible region. However, even bilinear coupling constraints in two-player games are generally nonconvex. Playerwise concavity, a relaxation of joint concavity, is a common structural assumption on utilities that ensures the existence of Nash equilibria in games without coupling constraints (Nikaidô and Isoda,, 1955). For games with coupling constraints where the feasible region is defined by shared, playerwise concave constraint functions, past work has not addressed the existence of Nash equilibria.
On the computational side, it is important to acknowledge two fundamental limitations: (a) Even for shared bilinear constraints, finding a feasible strategy is NP-hard (Witsenhausen,, 1986). And (b), even in games without coupling constraints, for a general-sum utility structure, finding a Nash equilibrium is PPAD-hard (Daskalakis et al.,, 2009).
To circumvent the intractability associated with (a), given an initial feasible strategy, one can ask whether there exist optimization methods under which players converge to a constrained Nash equilibrium. Under this initial feasibility assumption, recent work has provided positive answers in the context of constrained Markov potential games (Alatur et al.,, 2024; Jordan et al.,, 2024). Despite the limits posed by (b), convergent learning methods have been studied for several subclasses where Nash equilibria can be efficiently computed, including zero-sum games (Cesa-Bianchi and Lugosi,, 2006) and potential games (Heliou et al.,, 2017; Anagnostides et al.,, 2022).
To further position our work within the literature, we note that in the context of the variational inequality framework, a substantial body of work addresses games with coupling constraints. Recent results (Jordan et al.,, 2023; Tatarenko and Kamgarpour,, 2024) provide algorithms with convergence to constrained Nash equilibria for the class of monotone games with jointly linear constraints. Thus, their assumptions on utilities as well as constraints are stronger than our playerwise concavity. Coupling constraints that induce nonconvex feasible regions are considered in Scarabaggio et al., (2024), who do not assume playerwise concavity of constraints and characterize a weaker notion of local equilibria.
In summary, to the best of our knowledge, the class of continuous constrained games with playerwise concave objectives and shared, playerwise convex coupling constraints has not been explicitly studied. Applying prior results to this setting requires introducing restrictive assumptions. In this work, we focus on the existence and computation of equilibria in this class of games.
In the first part, we investigate the existence of constrained Nash equilibria. Our approach generalizes the result of Rosen, (1965) by relaxing the assumption on the feasible region from joint convexity to playerwise convexity. We build on tools from topological fixed point theory, specifically the works of Eilenberg and Montgomery, (1946) and Begle, (1950), which extend fixed point results from convex to more general, contractible domains. Analyzing the topological structure of the feasible region for the considered game class is a key technical contribution of our work.
In the second part of the paper, we address the computational problem. Our focus is on independent learning algorithms, where players simultaneously update their strategies using only gradient feedback on their own utility and the shared constraints, without requiring additional communication. We draw inspiration from log barrier regularized interior point methods. Originally proposed by Frisch, (1955), this approach has been extensively studied in the context of constrained optimization problems (Megiddo,, 1989; Hinder and Ye,, 2024), but has remained unexplored in the context of games. By carefully selecting adaptive learning rates, we guarantee convergence to approximate constrained Nash equilibria in a subclass of playerwise concave games with shared coupling constraints, namely when utilities admit a potential function.
@@
In summary, we consider continuous games with playerwise concave utilities and shared, playerwise concave coupling constraints. Our contributions are as follows:
  • We establish the existence of constrained Nash equilibria without relying on restrictive assumptions made in prior work. Our approach provides new insights into the topological structure of the feasible region.
  • We propose an independent log barrier regularized gradient method. Starting from an initial feasible strategy, players update their strategies based on exact gradient feedback for utility and constraint functions, using adaptive learning rates. For constrained potential games, we prove best-iterate convergence to a constrained $\epsilon$-approximate Nash equilibrium strategy at an $\mathcal{O}(\epsilon^{-3})$ rate.
  • We empirically demonstrate convergence of the proposed algorithm in an identical interest game with nonconvex feasible region, and in a network routing game with coupled link capacity constraints.
@@ 2 PRELIMINARIES
In this paper, we study static games with continuous strategy spaces. Let $[m]$ be the set of players. For each $i\in[m]$, $\mathcal{X}_{i}\subset\mathbb{R}^{d_{i}}$ is a compact convex set of available strategies. Joint strategy profiles are denoted by $x=(x_{1},\dots,x_{m})\in\mathcal{X}\coloneqq\times_{i\in[m]}\mathcal{X}_{i}$. Upon playing strategy $x\in\mathcal{X}$, each player $i$ receives utility $u_{i}(x)$ where $u_{i}:\mathcal{X}\to\mathbb{R}$ is continuous. Shared coupling constraints are introduced through the set of functions and thresholds $\{(c_{j},\alpha_{j})\}_{j\in[b]}$ where for each $j\in[b]$, $\alpha_{j}\in\mathbb{R}$, and $c_{j}:\mathcal{X}\to\mathbb{R}$ is a continuous function. Players are assumed to be utility maximisers subject to lower bounds on the constraint functions. The set of strategy profiles considered feasible is given by
@@
For joint strategies of all players other than $i$, we use the notation $x_{-i}\coloneqq(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{m})\in\mathcal{X}_{-i}\coloneqq\times_{j\in[m]\setminus\{i\}}\mathcal{X}_{j}$. For $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, let $\mathcal{C}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}\}$.
@@
Remark 1.
The term constrained game has been used in the literature with varying meanings. In some cases, it has been used to distinguish games over convex compact strategy spaces from unconstrained games over $\mathbb{R}^{n}$. Our notion of constrained game is more general, as it additionally allows for shared coupling constraints.
@@ 3 EXISTENCE OF CONSTRAINED NASH EQUILIBRIA
In this section, we establish our existence result for constrained Nash equilibria. We first discuss assumptions made in prior work, then give an overview of our approach.
3.1 Discussion of assumptions
In the following, we motivate our assumptions using two examples of constrained games. Example 1 serves to illustrate that Assumption 1 is not sufficient for the existence of constrained Nash equilibria as it allows for feasible sets with nonsmooth boundary. We then introduce Assumption 2 to prevent the occurence of such degeneracies. Finally, we present Example 2, which satisfies both Assumption 1 and Assumption 2. At the same time, Example 2 illustrates that these assumptions are weaker than the conditions imposed by earlier existence results in the literature.
@@ 3.1 Discussion of assumptions
Consider a two-player game with continuous strategy spaces $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and feasible set $\mathcal{C}^{(1)}$ defined by a constraint $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$ with threshold $\alpha_{1}=0$; illustrated on the left-hand side of Figure 1. Moreover, define the players’ utilities as $u_{1}(x_{1},x_{2})=u_{2}(x_{1},x_{2})\coloneqq x_{1}(1-x_{2})$. Observe that no strategy $(x_{1},x_{2})\in\mathcal{C}^{(1)}$ with $x_{1}<\frac{1}{2}$ can be a constrained Nash equilibrium, as $u_{1}$ would increase if player 1 would switch to $x_{1}^{\prime}=\frac{1}{2}$ and since $(x_{1}^{\prime},x_{2})\in\mathcal{C}^{(1)}$. By an analogous argument, no $(x_{1},x_{2})\in\mathcal{C}^{(1)}$ with $x_{2}>\frac{1}{2}$ can be a constrained Nash equilibrium. The only point not covered by above two cases is $(\frac{1}{2},\frac{1}{2})$ which however is not a Nash equilibrium either.
We note that in Example 1, the gradient of the constraint $c_{1}$ vanishes at a point that lies on the boundary of $\mathcal{C}^{(1)}$, namely $(\frac{1}{2},\frac{1}{2})$. In order to prove our existence result, we need to rule out such degeneracies. We therefore require the following playerwise variant of the Mangasarian-Fromovitz constraint qualification condition (Mangasarian and Fromovitz,, 1967); a regularity assumption that guarantees a uniform descent direction for all constraints at points on the boundary of $\mathcal{C}$. In our proof, this is used argue about continuity of best response mappings.
@@
Finally, we present an example of a constrained game that satisfies Assumptions 1 and 2. Moreover, its feasible region $\mathcal{C}^{(2)}$ does not satisfy the stronger conditions required by previous existence results.
@@
Let again $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and consider the feasible set $\mathcal{C}^{(2)}$ defined by two constraints, $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})-\frac{1}{111}$ and $c_{2}(x_{1},x_{2})\coloneqq\frac{1}{11}-(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$, with thresholds $\alpha_{1}=\alpha_{2}=0$, as shown on the right-hand side of Figure 2. The utility structure is irrelevant for this example. The resulting feasible set $\mathcal{C}^{(2)}$ is illustrated as the blue region on the right-hand side of Figure 1. The constraints satisfy Assumption 1, as fixing $x_{1}$ yields a linear function in $x_{2}$, and vice versa. Moreover, Assumption 2 holds, as $|\mathcal{J}(x)|=1$ for all $x\in\mathcal{C}^{(2)}$ and neither $c_{1}$ nor $c_{2}$ has any critical points on the boundary of $\mathcal{C}^{(2)}$. One can also observe that $\mathcal{C}^{(2)}$ is not convex, hence Rosen, (1965)’s result is not applicable. Moreover, for any $x\in\mathcal{X}$ with $x_{1}=0.5$, we have $\mathcal{C}^{(2)}_{2}(x_{1})=\emptyset$. Hence, results relying on the nonemptiness of feasible responses such as Debreu, (1952) are not applicable either.
Feasible regions similar to the one in Example 2 naturally arise in matrix games, where nonconvexity stems from joint probability products, and the nonemptiness condition fails when satisfying the constraint requires some coordination among players to choose certain feasible joint actions, see e.g. the applications in Section 5.
@@ 3.2 Existence result
We now state our existence theorem. Note that clearly no Nash equilibrium can exist when $\mathcal{C}=\emptyset$. {restatable}theoremthmexistence Let Assumptions 1 and 2 hold and suppose $\mathcal{C}\not$$\emptyset$. Then there exists a constrained Nash equilibrium $x^{\star}\in\mathcal{X}$. The remainder of this section is devoted to giving an overview of our proof approach and introducing the relevant topological background.
Fixed point theory
Existence proofs for Nash equilibria typically proceed by constructing a best response map to which a fixed point theorem is applied. Classical results such as Brouwer’s or Kakutani’s theorem require the domain of this map to be convex. In order to establish the existence of a fixed point for a best response map defined over our nonconvex $\mathcal{C}$, we rely on a result that goes back to Eilenberg and Montgomery, (1946), and Begle, (1950), which relaxes the convexity requirement to the weaker topological notion of contractibility. Below, we define contractibility and state a version of Begle, (1950)’s result applied to finite-dimensional Euclidean space. For examples of contractible sets, and a proof of Fact 3.2, see Appendix C. {restatable}definitiondefcontractibility Let $S\subseteq\mathbb{R}^{n}$. We say $S$ is contractible if there exists $s_{0}\in S$, and a continuous map $H:[0,1]\times S\to S$ such that for all $s\in S$, it holds that $H(0,s)=s$ and $H(1,s)=s_{0}$. {restatable}factfactfp Let $X\subset\mathbb{R}^{n}$ be compact, contractible, and $\phi:X\to X$ continuous. Then there exists $x\in X$ such that $x=\phi(x)$, that is, $x$ is a fixed point of $\phi$.
As seen in Example 2, the feasible region may be disconnected; thus, it cannot be contractible. Therefore, we define a best response function only over a connected component of $\mathcal{C}$, to which we then apply Fact 3.2. We conclude this section by sketching the proof of Theorem 3.2. The formal proof is given in Appendix D.
Figure 2: Proof sketch for contractability of $\mathcal{C}^{\prime}$.
Proof overview
Let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be any connected component of $\mathcal{C}$, which exists as $\mathcal{C}\not$$\emptyset$. For $x_{-i}\in\mathcal{X}_{-i}$, we use the notation $\mathcal{C}^{\prime}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}^{\prime}\}$ and $\operatorname{supp}(\mathcal{C}_{i}^{\prime})\coloneqq\left\{x_{-i}\in\mathcal{X}_{-i}\mid\mathcal{C}^{\prime}_{i}(x_{-i})\not$$\emptyset\right\}$. The proof is divided into two parts:
  1. 1)
    Contractibility of $\mathcal{C}^{\prime}$: We construct a mapping $H:[0,1]\times\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ by sequentially contracting the feasible region along the dimensions of the players. For this overview, we focus on the two-player case which is illustrated in Figure 2.
@@
    Property (a) is shown via convexity of the slices $\mathcal{C}^{\prime}_{i}(x_{-i})$ which follows directly from the playerwise concavity on the constraints. Regarding (b), we observe that $\overline{c}(\cdot)$ and $\mathcal{C}_{i}^{\prime}(\cdot)$, from which we compose $H$, are both continuous. In particular, continuity of the set-valued mapping $\mathcal{C}_{i}^{\prime}(\cdot)$ follows from standard results in parametric optimization, which apply under the MFCQ condition (Assumption 2). For (c), by construction, for all $x\in\mathcal{C}^{\prime}$, $H(0,x)=x$ and $H(1,x)=\overline{c}(\operatorname{supp}\mathcal{C}_{1}^{\prime})$.
  1. 2)
    Fixed point argument: Our goal is to define a best response function $\Psi:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$, to which Fact 3.2 can be applied. We leverage the playerwise concave constraint structure by constructing $\Psi$ in a playerwise sequential manner. For each $i\in[m]$, suppose $u_{i}$ is strongly concave (we will justify why this is without loss of generality). Then $\Psi_{i}:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ with
@@
    is well-defined since the maximizer is unique. Moreover, by a variant of Berge’s maximum theorem for strongly concave objective functions over convex constraints (see Theorem 1 in Appendix C), $\Psi_{i}$ is continuous. We then apply Fact 3.2 to the composition $\Psi\coloneqq\Psi_{1}\circ\dots\circ\Psi_{m}$. Using the playerwise concavity of constraints, it can be argued that any fixed point $x^{\star}\in\mathcal{C}^{\prime}$ with $x^{\star}=\Psi(x^{\star})$ is a constrained Nash equilibrium.
    Why is assuming strict concavity of $u_{i}$ without loss of generality? For $\epsilon>0$, consider the perturbed game with playerwise strongly concave utilities $u_{i}^{\epsilon}(x)\coloneqq u_{i}(x)-\epsilon\lVert x_{i}\rVert^{2}$. By a limit argument, we can show that for a sequence of such games with $\epsilon\to 0$, any limit point of the respective constrained Nash equilibria (just shown to exist) is a constrained Nash equilibrium of the original unperturbed game.
    We point out that, different from other existence proofs (Debreu,, 1952; Rosen,, 1965), it is important for the best response maps $\Psi_{i}$ to be single-valued: Otherwise it is not clear how to ensure $\Psi$ to be convex-valued (or contractible-valued), which is necessary for fixed point theorems to hold.
Having established existence, we next investigate settings in which such equilibria can be computed or learned in an efficient manner.

### v2

@@ 1 INTRODUCTION
Optimization problems involving multiple self-interested agents with coupled objectives frequently arise in applications across robotics, computer science, and economics. While in standard game theoretic frameworks, the players’ feasible action sets are decoupled, in many interactive tasks, players are restricted to optimizing over a set of admissible strategies that are influenced by the choices of others. This game setting is captured by introducing coupling constraints, that is, constraints which depend on the players’ joint strategy profile. For instance, in economic modeling, such constraint games have been applied to analyze equilibria under resource limitations and market constraints (Arrow and Debreu,, 1954). In other applications, players may be subject to joint safety requirements such as collision avoidance in multi-robot control (Gu et al.,, 2023), or operating within a shared budget, such as limited bandwidth in communication networks (Lasaulce and Tembine,, 2011). Despite their high relevance, several questions about the existence and computation of equilibria in this class of games remain open. In this paper, we address some of these questions, focusing on concave games with playerwise concave coupling constraints. To clarify our contributions, we first discuss the state of the art on games with coupling constraints.
Existence
@@
Shortly after Nash’s foundational work on existence of mixed strategy equilibria in noncooperative finite action games, Debreu (Debreu,, 1952) introduced games with coupling constraints. In his formulation, each player may face a coupling constraint distinct from those of other players—a setting we refer to as non-shared coupling constraints. The equilibrium of such a game, called a social equilibrium and later referred to as a constrained Nash equilibrium, is one where no player has an incentive to unilaterally deviate within her set of feasible actions. Debreu showed that such equilibria exist in games with compact, convex strategy spaces and playerwise quasi-concave utilities, provided that each player’s set of feasible responses remains nonempty, irrespective of the strategies chosen by others. This requirement is often restrictive: in many practical scenarios, such as under shared safety constraints, violations caused by a subset of players may leave others unable to restore feasibility. However, this assumption appears necessary in Debreu’s setting, as without it one can construct examples—even with linear utilities and non-shared linear coupling constraints—where no constrained Nash equilibrium exists (see Appendix B).
Later, Rosen, (1965) showed that for the class of constrained games in which all players have concave utilities and share the same coupling constraints, constrained Nash equilibria exist without requiring the nonemptiness of feasible response sets as in Debreu, (1952). However, this work assumed a jointly convex feasible region. This assumption is restrictive, since even bilinear coupling constraints in two-player settings are generally nonconvex. Such a bilinear structure arises naturally for instance in mixed strategy extensions of finite action games. Thus, an open question is whether existence of Nash equilibria can be guaranteed in games with shared coupling constraints under conditions weaker than joint convexity or nonemptiness of feasible response sets?
Playerwise concavity—a relaxation of joint concavity—is a common structural assumption on utilities that ensures the existence of Nash equilibria in games without constraints (Rosen,, 1965). We consider the natural extension of this game class to settings with coupling constraints and prove existence of constrained Nash equilibria for games with playerwise concave utilities and playerwise concave coupling constraints.
Computation
While existence results are fundamental, for Nash equilibria to be plausible outcomes of multi-agent interactions, they should also be computationally tractable. It is known that adding constraints to games can introduce significant computational challenges. For instance, it was recently shown that convex-concave zero-sum two-player games with coupling constraints are computationally intractable (Daskalakis et al.,, 2021; Bernasconi et al.,, 2024), a stark contrast to their unconstrained counterparts.
Similar to prior existence results, much of the literature on computing constrained Nash equilibria relies on restrictive assumptions about the game structure. For example, in monotone games with a jointly convex feasible region, coupling constraints can be handled using the variational inequality framework (Facchinei and Pang,, 2003). These assumptions on both utilities and constraints are stronger than playerwise concavity.
Unfortunately, even for shared bilinear constraints, finding a feasible strategy is NP-hard (Witsenhausen,, 1986). However, when an initial feasible strategy is given, one can ask whether there exist learning dynamics—understood here as iterative update procedures—under which players converge to a constrained Nash equilibrium. Under this initial feasibility assumption, recent work has provided positive answers in the context of constrained Markov potential games (Alatur et al.,, 2024; Jordan et al.,, 2024), a setting with finite action spaces. We provide such a convergent learning method for the class of continuous action potential games with playerwise concave utilities and shared playerwise concave coupling constraints.
@@
In summary, our contributions are as follows.
  1. 1.
    For the class of games with playerwise convex feasible regions, we establish the existence of constrained Nash equilibria without requiring nonemptiness of feasible response sets. Our approach generalizes the result of Rosen, (1965) by relaxing the assumption on the feasible region from joint convexity to playerwise convexity.
    Insights. We build on tools from topological fixed point theory, specifically the works of Eilenberg and Montgomery, (1946), and Begle, (1950), which extend fixed point results from convex to more general, contractible domains. A key technical contribution of our work is the analysis of the topological structure of the feasible region, where we show that playerwise concavity of constraints yields a contractibility property.
  2. 2.
    Restricting the above class to potential games with coupling constraints, we develop an independent learning algorithm with provable convergence to a constrained Nash equilibrium at an $\mathcal{O}(\epsilon^{-3})$ rate. Players simultaneously update their strategies using only gradient feedback on their own utility and the shared constraints, without requiring additional communication.
    Insights. We draw inspiration from log barrier regularized interior point methods. Originally proposed by Frisch, (1955), this approach has been extensively studied in the context of constrained optimization problems (Megiddo,, 1989; Hinder and Ye,, 2024), but has remained unexplored in the context of games. Starting from an initial feasible strategy, and by carefully selecting adaptive learning rates, we guarantee convergence along the potential function to approximate equilibria.
  3. 3.
    We empirically demonstrate convergence of the proposed algorithm in a cooperative game with a nonconvex feasible region and in a network routing game with coupled link capacity constraints.
Paper outline
We begin by introducing the constrained game setting in Section 2. In Section 3, we provide our existence result. Section 4 then focuses on learning dynamics in constrained potential games. Lastly, we describe our simulations in Section 5 and conclude the paper with Section 6.
@@ 2 PRELIMINARIES
In this paper, we study static games with continuous strategy spaces. Let $[m]$ be the set of players. For each $i\in[m]$, $\mathcal{X}_{i}\subset\mathbb{R}^{d_{i}}$ is a compact convex set of available strategies. Joint strategy profiles are denoted by $x=(x_{1},\dots,x_{m})\in\mathcal{X}\coloneqq\times_{i\in[m]}\mathcal{X}_{i}$. Upon playing strategy $x\in\mathcal{X}$, each player $i$ receives utility $u_{i}(x)$ where $u_{i}:\mathcal{X}\to\mathbb{R}$ is continuous. Shared coupling constraints are introduced through a set of functions and thresholds $\{(c_{j},\alpha_{j})\}_{j\in[b]}$ where for each $j\in[b]$, $\alpha_{j}\in\mathbb{R}$, and $c_{j}:\mathcal{X}\to\mathbb{R}$ is a continuous function. Players are assumed to be utility maximisers subject to lower bounds on the constraint functions. The set of strategy profiles considered feasible is given by
@@
For joint strategies of all players other than $i$, we use the notation $x_{-i}\coloneqq(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{m})\in\mathcal{X}_{-i}\coloneqq\times_{j\in[m]\setminus\{i\}}\mathcal{X}_{j}$. For $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, let $\mathcal{C}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}\}$ denote the feasible response set of player $i$.
@@ 3 EXISTENCE OF CONSTRAINED NASH EQUILIBRIA
In this section, we establish our existence result for constrained Nash equilibria. We begin by discussing the structure of the feasible region induced by playerwise concave constraints. We then identify minimal conditions that guarantee existence of equilibria and illustrate how these results generalize prior work. Finally, we give an overview of our proof approach.
3.1 Structure of the feasible region
Existence proofs for Nash equilibria typically proceed by applying a fixed point theorem to a continuous best response function. In order to define this best response map over players’ feasible response sets, for each player $i$, $\mathcal{C}_{i}(x_{-i})$ should vary continuously as a set-valued function of $x_{-i}$. The following example demonstrates that playerwise concavity of constraints is not sufficient for this continuity property to hold. Moreover, Example 1 shows that constrained Nash equilibria may not exist in games with feasible regions that lack such continuity of the feasible response map.
@@ 3.1 Structure of the feasible region
Consider a two-player game with continuous strategy spaces $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and feasible set $\mathcal{C}^{(1)}$ defined by a constraint $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$ with threshold $\alpha_{1}=0$; illustrated on the left-hand side of Figure 1. Moreover, define the players’ utilities as $u_{1}(x_{1},x_{2})=u_{2}(x_{1},x_{2})\coloneqq x_{1}(1-x_{2})$. Observe that no strategy $(x_{1},x_{2})\in\mathcal{C}^{(1)}$ with $x_{1}<\frac{1}{2}$ can be a constrained Nash equilibrium, since then for $x_{1}^{\prime}=\frac{1}{2}$, we have $u_{1}(x_{1}^{\prime},x_{2})>u_{1}(x_{1},x_{2})$ and $(x_{1}^{\prime},x_{2})\in\mathcal{C}^{(1)}$. By an analogous argument, no $(x_{1},x_{2})\in\mathcal{C}^{(1)}$ with $x_{2}>\frac{1}{2}$ can be a constrained Nash equilibrium. The only point not covered by above two cases is $(\frac{1}{2},\frac{1}{2})$ which however is not a Nash equilibrium either.
We note that in Example 1, the discontinuity in the set-valued mapping $x_{1}\mapsto\mathcal{C}_{2}(x_{1})$ occurs at $x_{1}=\frac{1}{2}$ where the map is upper but not lower-semicontinuous (see Appendix C for definitions). In order to rule out such degeneracies, we require that for all points on the boundary of the feasible region $\mathcal{C}$, there exists a uniform descent direction for all constraints. Together with continuity of the constraints $c_{j}(\cdot)$, this ensures that for any point on the boundary of $\mathcal{C}$ and any player $i$, there exists a neighborhood of feasible responses for player $i$ that is stable under small perturbations of the other players’ strategies.
@@
We point out that this condition is a playerwise variant of the Mangasarian-Fromovitz constraint qualification condition (Mangasarian and Fromovitz,, 1967).
Next, in order to demonstrate the generality of our conditions, we present an example of a constrained game whose feasible region $\mathcal{C}^{(2)}$ satisfies our Assumptions 1 and 2, but does not fulfill the joint convexity or nonempty feasible response requirements imposed by previous existence results. Feasible regions similar to the one in Example 2 naturally arise in matrix games, where nonconvexity stems from joint probability products, and the nonemptiness condition fails when satisfying the constraint requires some coordination among players to choose certain feasible joint actions.
@@
Let again $\mathcal{X}_{1}=\mathcal{X}_{2}=[0,1]$ and consider the feasible set $\mathcal{C}^{(2)}$ defined by two constraints, $c_{1}(x_{1},x_{2})\coloneqq(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$ and $c_{2}(x_{1},x_{2})\coloneqq-(x_{1}-\frac{1}{2})(x_{2}-\frac{1}{2})$, together with thresholds fixed, for concreteness, at $\alpha_{1}=\frac{1}{111},\alpha_{2}=-\frac{1}{11}$, as shown on the right-hand side of Figure 2. The utility structure is irrelevant for this example. The resulting feasible set $\mathcal{C}^{(2)}$ is illustrated as the blue region on the right-hand side of Figure 1. The constraints satisfy Assumption 1, as fixing $x_{1}$ yields a linear function in $x_{2}$, and vice versa. Moreover, Assumption 2 holds, as $|\mathcal{J}(x)|=1$ for all $x\in\mathcal{C}^{(2)}$ and neither $c_{1}$ nor $c_{2}$ has any critical points on the boundary of $\mathcal{C}^{(2)}$. One can also observe that $\mathcal{C}^{(2)}$ is not convex, hence Rosen, (1965)’s result is not applicable. Moreover, for any $x\in\mathcal{X}$ with $x_{1}=0.5$, we have $\mathcal{C}^{(2)}_{2}(x_{1})=\emptyset$. Hence, results relying on the nonemptiness of feasible responses such as Debreu, (1952) are not applicable either.
In the following, we address the existence of Nash equilibria for games with feasible regions such as in Example 2, namely the class of constrained games captured by Assumptions 1 and 2.
@@ 3.2 Existence result
{restatable}
theoremthmexistence Let Assumptions 1 and 2 hold and suppose $\mathcal{C}\not$$\emptyset$. Then there exists a constrained Nash equilibrium $x^{\star}\in\mathcal{X}$. The full proof of Theorem 3.2 is given in Appendix D; here we sketch the main elements.
A key challenge in our setting is the nonconvexity of the feasible region $\mathcal{C}$. Fixed point theorems such as Brouwer’s or Kakutani’s, which are typically used in existence proofs for Nash equilibria, require the feasible set to be convex. In order to establish the existence of a fixed point for a best response map defined over our nonconvex $\mathcal{C}$, we rely on a result that goes back to Eilenberg and Montgomery, (1946), and Begle, (1950), which relaxes the convexity requirement to the weaker topological notion of contractibility. A central part of our proof, Lemma 3.2, establishes that under Assumption 2, playerwise concavity of constraints implies contractibility of each connected component of $\mathcal{C}$.
Below, we define contractibility and state a version of Begle, (1950)’s theorem applied to finite-dimensional Euclidean space. For examples of contractible sets, and a proof of Fact 3.2, see Appendix C. {restatable}definitiondefcontractibility Let $S\subseteq\mathbb{R}^{n}$. We say $S$ is contractible if there exists $s_{0}\in S$, and a continuous map $H:[0,1]\times S\to S$ such that for all $s\in S$, it holds that $H(0,s)=s$ and $H(1,s)=s_{0}$. {restatable}factfactfp Let $X\subset\mathbb{R}^{n}$ be compact, contractible, and $\phi:X\to X$ continuous. Then there exists $x\in X$ such that $x=\phi(x)$, that is, $x$ is a fixed point of $\phi$.
As seen in Example 2, the feasible region $\mathcal{C}$ may be disconnected; thus, it cannot be contractible. However, we establish the following key contractibility result.
{restatable}
lemmalemcontract Suppose Assumptions 1 and 2 hold. Let $\mathcal{C}\not$$\emptyset$ and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be any connected component of $\mathcal{C}$. Then $\mathcal{C}^{\prime}$ is contractible.
Figure 2: Proof sketch for contractibility (Lemma 3.2).
Proof overview for Lemma 3.2
Because $\mathcal{C}^{\prime}$ is not convex, we cannot simply contract it to a point via linear interpolation within $\mathcal{C}^{\prime}$. However, we may leverage the playerwise concavity structure and construct a mapping $H:[0,1]\times\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ by sequentially contracting the feasible region along the dimensions of the players. For this overview, we focus on the two-player case which is illustrated in Figure 2. For $x_{-i}\in\mathcal{X}_{-i}$, let $\mathcal{C}^{\prime}_{i}(x_{-i})\coloneqq\{x_{i}\in\mathcal{X}_{i}\mid(x_{i},x_{-i})\in\mathcal{C}^{\prime}\}$ and $\operatorname{supp}(\mathcal{C}_{i}^{\prime})\coloneqq\left\{x_{-i}\in\mathcal{X}_{-i}\mid\mathcal{C}^{\prime}_{i}(x_{-i})\not$$\emptyset\right\}$.
@@
Property (a) is shown via convexity of the slices $\mathcal{C}^{\prime}_{i}(x_{-i})$ which follows directly from the playerwise concavity of the constraints. Regarding (b), we observe that $\overline{c}(\cdot)$ and $\mathcal{C}_{i}^{\prime}(\cdot)$, from which we compose $H$, are both continuous. In particular, continuity of the set-valued mapping $\mathcal{C}_{i}^{\prime}(\cdot)$ follows from standard results in parametric optimization, which apply under the MFCQ condition (Assumption 2). For (c), by construction, for all $x\in\mathcal{C}^{\prime}$, $H(0,x)=x$ and $H(1,x)=\overline{c}(\operatorname{supp}(\mathcal{C}_{1}^{\prime}))$.
Having established contractibility of $\mathcal{C}^{\prime}$, we prove Theorem 3.2 via a fixed point argument.
Proof overview for Theorem 3.2
Our goal is to define a best response function $\Psi:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$, to which Fact 3.2 can be applied. The usual construction of simultaneously mapping players’ strategies to their best responses does not work here, as due to the nonconvexity of $\mathcal{C}^{\prime}$, the resulting strategies may not be feasible. Instead, we again leverage the playerwise concave constraint structure by constructing $\Psi$ in a playerwise sequential manner. For each $i\in[m]$, suppose $u_{i}$ is strongly concave (see limit argument in the full proof, Appendix D, for why this is without loss of generality). Then $\Psi_{i}:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ with
@@
is well-defined since the maximizer is unique. Moreover, by a variant of Berge’s maximum theorem for strongly concave objective functions over convex constraints (see Appendix C), $\Psi_{i}$ is continuous. We then apply Fact 3.2 to the composition $\Psi\coloneqq\Psi_{1}\circ\dots\circ\Psi_{m}$. Using the playerwise concavity of constraints, it can be argued that any fixed point $x^{\star}\in\mathcal{C}^{\prime}$ with $x^{\star}=\Psi(x^{\star})$ is a constrained Nash equilibrium.